# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Video

### v1

@@ Methods
Restormer [92] 33.33 / 0.974 / 0.043 32.43 / 0.962 / 0.064
DSTNet [56] 33.13 / 0.971 / 0.039 32.42 / 0.962 / 0.059
BSSTNet [95] 34.13 / 0.980 / 0.027 33.63 / 0.973 / 0.039
DeformableGS [90] 30.92 / 0.949 / 0.103 29.11 / 0.926 / 0.191
4DGaussians [84] 31.25 / 0.949 / 0.097 29.95 / 0.932 / 0.169
E-D3DGS [1] 31.28 / 0.949 / 0.098 29.98 / 0.931 / 0.170
DyBluRF [72] 29.87 / 0.947 / 0.054 29.23 / 0.943 / 0.105
Shape-of-Motion [79] 29.83 / 0.946 / 0.107 28.93 / 0.932 / 0.190
Deblur4DGS (Ours) 32.55 / 0.975 / 0.041 31.60 / 0.965 / 0.084
5.2 Comparison with State-of-the-Art Methods
We compare Deblur4DGS with 5 state-of-the-art 4D reconstruction methods (i.e., DeformableGS [90], 4DGaussians [84], E-D3DGS [1], DyBluRF [72], and Shape-of-Motion [79]) on two resolution (i.e., $228\times 512$ and $720\times 1080$) datasets. We input poses of right and left cameras to produce novel-view synthesis and deblurring results, respectively. We linearly interpolate the left camera poses and timestamps to generate ×$\times 16$ frame interpolation results. We employ a Gaussian filter to smooth left camera poses for video stabilization, following  [59]. The results for novel-view synthesis, deblurring, frame interpolation, and video stabilization are summarized in Tab. A, Tab. B, Tab. 4, and Tab. 5, respectively. Besides, rendering speed comparisons on $720\times 1080$ images are reported in Fig. 1.
Novel-view synthesis Comparisons. As shown in Tab. A, among all competing methods, DyBluRF [72] gets the best LPIPS score, while Shape-of-Motion [79] achieves the best PSNR and SSIM scores. Benefiting the explicit 3D representation manner and elaborate design for processing motion blur, our Deblur4DGS obtains all the best scores on both datasets. In addition, Deblur4DGS removes blur more clearly and produces more photo-realistic details in both static and dynamic areas, as shown in Fig. 4. More visual comparisons can be seen in the Suppl.
In addition, to further demonstrate the effectiveness of Deblur4DGS, we first pre-process the blurry videos with state-of-the-art image [92] or video [56, 95] deblurring methods and then perform 4D reconstruction. The results are summarized in Tab. 2. Compared with reconstruction from blurry videos, the incorporation of deblurring models improves performance. This is because the deblurring models remove some blur, facilitating sharp scene reconstruction. However, as the image and video deblurring methods cannot perceive 3D structure and maintain scene geometric consistency, the scene reconstruction results are still unsatisfactory. In contrast, Deblur4DGS jointly reconstructs scene geometry and processes motion blur in 3D space, achieving better scene reconstruction results.
Table 4: Quantitative comparisons for frame interpolation. The first category is supervised video frame interpolation methods, and the second is 4D reconstruction-based ones.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
RIFE [21] 0.179 / 44.090 0.147 / 32.637
EMAVFI [94] 0.181 / 42.773 0.170 / 32.564
VIDUE [66] 0.267 / 60.854 0.185 / 49.560
DeformableGS [90] 0.172 / 41.603 0.167 / 31.205
4DGaussians [84] 0.177 / 44.132 0.177 / 33.836
E-D3DGS [1] 0.178 / 43.250 0.181 / 32.945
DyBluRF [72] 0.166 / 51.513 0.132 / 36.768
Shape-of-Motion [79] 0.176 / 43.020 0.202 / 33.817
Deblur4DGS (Ours) 0.200 / 54.043 0.213 / 42.139
Table 5: Quantitative comparisons for video stabilization. The first category is 2D video stabilization methods, and the second is 4D reconstruction-based ones.
Methods
[…]
$288\times 512$
[…]
$720\times 1080$
MeshFlow [47] 0.140 / 38.289 0.116 / 34.140
NNDVS [100] 0.186 / 45.613 0.137 / 33.269
DeformableGS [90] 0.170 / 40.846 0.168 / 30.971
4DGaussians [84] 0.145 / 44.673 0.172 / 34.416
E-D3DGS [1] 0.181 / 44.364 0.182 / 33.828
DyBluRF [72] 0.164 / 53.115 0.132 / 37.556
Shape-of-Motion [79] 0.174 / 43.293 0.202 / 33.750
Deblur4DGS (Ours) 0.199 / 54.299 0.211 / 42.241
Deblurring Comparisons. Apart from 4D reconstruction-based methods, we additionally compare with some state-of-the-art image [92] and video [56, 95] deblurring ones. As shown in Tab. B, Deblur4DGS obtains better results than 4D reconstruction-based methods and comparable ones to the deblurring-specific (i.e., Restormer [92], DSTNet [56], and BSSTNet [95]) methods. Compared with the former, Deblur4DGS better reconstructs the scene, thus performing better. Note that the latter ones are trained on large paired data in a fully supervised manner while Deblur4DGS is optimized only with the given blurry video in a self-supervised manner. Although the large data prior makes these methods perform better, Deblur4DGS is more convenient to use without paired data collection.
Frame Interpolation Comparisons. We compare Deblur4DGS with 4D reconstruction-based methods and some state-of-the-art supervised video frame interpolation ones (i.e., RIFE [21], EMAVFI [94], and VIDUE [66]), as shown in Tab. 4. Deform4DGS generally performs satisfactorily. Note that VIDUE [66] is specially designed and trained with large paired data in a supervised manner for joint deblurring and fame Interpolation, thus achieving better results.
Video Stabilization Comparisons. The results are summarized in Tab. 5. Benefiting from the better geometry reconstruction, Deblur4DGS achieves pleasant scores compared with both 2D video stabilization methods (i.e., MeshFlow [47] and NNDVS [100]) and 4D reconstruction-based ones, which further demonstrates its effectiveness.
Figure 5: Effect about the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG). Static areas are marked with yellow boxes while dynamic ones are marked with red boxes. Without ECP and EDG, the results are almost as blurry as the input frame, as shown in (b). Only with ECP, static areas are sharp but may lead to visual artifacts in dynamic areas, as shown in (c). Further introducing EDG improves this, as shown in (d).
6 Ablation Study
We first discuss the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG). Then, we validate the effect of regularization terms, as well as the blur-aware variable (BAV) canonical Gaussians.
6.1 Effect of ECP and EDG
ECP and EDG are introduced to process camera motion blur and object motion blur, respectively. Visual comparisons are shown in Fig. 5, and quantitative results are in the Suppl. First, without ECP and EDG, the results are almost as blurry as the input frame, as shown in Fig. 5(b). Second, only with ECP, the static areas are sharp but may lead to visual artifacts in dynamic areas, as shown in Fig. 5(c). It is because ECP cannot simulate the object movement. Third, we further introduce EDG to simulate that produces visually pleasant results in both areas, as shown in Fig. 5(d).
6.2 Effect of Regularization Terms
The effect of exposure regularization $\mathcal{L}_{e}$ (see Eq. 12, multi-frame consistency regularization $\mathcal{L}_{mfc}$ (see Eq. 13) and multi-resolution consistency regularization $\mathcal{L}_{mrc}$ (see Eq. 14) are shown in Tab. D. Without these regularization terms, the performance drops significantly and the visual results in Fig. 3 show that dynamic areas are with notable artifacts. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ alleviates the trivial solutions and improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. Their combinations perform best.
6.3 Effect of BAV Canonical Gaussians.
The effect of BAV canonical Gaussians is summarized in Tab. E. First, selecting a single canonical Gaussians across the entire video (i.e., None) leads to poor performance, due to the challenge of modeling large object motion. Second, selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates the challenge, thus leading to performance improvement. We also experiment with an optical flow-based strategy [67] to select the canonical Gaussians, and it achieves a similar performance to the uniform selection. This may be due to the inaccurate estimation of optical flow from blurry images. Third, our blur-aware selection is better, as the canonical Gaussians corresponding to the sharper frame help blur removal.
Table 6: Ablation studies on regularization terms (see Eq. 15).
[…]
$\times$ $\times$ $\times$ 28.69 / 0.947 / 0.063
$\times$ $\checkmark$ $\checkmark$ 28.86 / 0.949 / 0.060
$\checkmark$ $\times$ $\times$ 28.73 / 0.947 / 0.060
$\checkmark$ $\checkmark$ $\times$ 28.87 / 0.948 / 0.060
$\checkmark$ $\times$ $\checkmark$ 28.86 / 0.948 / 0.060
$\checkmark$ $\checkmark$ $\checkmark$ 28.92 / 0.949 / 0.060
Table 7: Effect of blur-aware variable (BAV) canonical Gaussians. ‘None’ denotes selecting a single one across the entire video.
[…]
None 28.67 / 0.946 / 0.065
w/o Blur-Aware 28.85 / 0.949 / 0.062
Ours 28.92 / 0.949 / 0.060
7 Conclusions
In this work, we propose Deblur4DGS, the first 4D Gaussian Splatting framework to reconstruct a high-quality 4D model from blurry monocular video. In particular, with the explicit motion trajectory modeling based on 3D Gaussian Splatting, we propose to transform the challenging continuous dynamic representation estimation within an exposure time into the exposure time estimation, where a series of regularizations are suggested to tackle the under-constrained optimization. Besides, a blur-aware variable canonical Gaussians is present to represent objects with large motion better. Beyond novel-view synthesis, Deblur4DGS can improve blurry video quality from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive results show Deblur4DGS outperforms state-of-the-art 4D reconstruction methods.
[…]
@@ 7 Conclusions
  • More result comparisons in Sec. C.
  • Effect of ECP and EDG in Sec. D.
  • Effect of regularization terms in Sec. E.
  • Effect of BAV canonical Gaussians in Sec. F.
@@ A Structure of Camera Motion Predictor
The structure of camera motion predictor is provided in Fig. A. It first embeds the camera pose $\mathbf{P}_{t}$ to a higher dimensional space, following [52]. Then, three FC blocks are stacked, each consisting of an FC layer followed by a ReLU operation. Finally, we deploy two heads to predict the camera pose at exposure start and end (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$), respectively.
@@ B Other Regularization Terms
Following Shape-of-Motion [79], we use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, including mask regularization $\mathcal{L}_{mask}$, 2D tracks regularization $\mathcal{L}_{track}$ and distance-preserving regularization $\mathcal{L}_{rigid}$.
@@
$\hat{\mathbf{M}}_{t}^{B}(u,v)=\texttt{max}\{\hat{\mathbf{M}}_{t,1}(u,v),\hat{\mathbf{M}}_{t,2}(u,v),...,\hat{\mathbf{M}}_{t,N}(u,v)\}.$ (A)
@@
$\mathcal{L}_{mask}=\mathcal{L}_{1}(\hat{\mathbf{M}}_{t}^{B},\mathbf{M}_{t}),$ (B)
where $\mathbf{M}_{t}$ is the mask that obtained by applying SAM2 [61] to the ground truth blurry frame. Besides, we render the 2D tracks $\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}$ from a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise it by the lifted long-range 2D tracks $\mathbf{U}_{t\rightarrow t^{\prime}}$ that are extracted from TAPIR [9], i.e.,
$\mathcal{L}_{track}=\mathcal{L}_{1}(\hat{\mathbf{U}}_{t\rightarrow t^{\prime}},\mathbf{U}_{t\rightarrow t^{\prime}}).$ (C)
@@
$\mathcal{L}_{rigid}=||\texttt{dist}(\hat{\mathbf{x}}_{t},\mathcal{C}_{J}(\mathbf{x}_{t}))-\texttt{dist}(\hat{\mathbf{x}}_{t^{\prime}},\mathcal{C}_{J}(\mathbf{x}_{t^{\prime}}))||_{2}^{2}.$ (D)
@@
Figure A: Structure of camera motion predictor.
@@
$\mathcal{L}_{oth}=\lambda_{mask}\mathcal{L}_{mask}+\lambda_{track}\mathcal{L}_{track}+\lambda_{rigid}\mathcal{L}_{rigid}.$ (E)
@@
C More Result Comparisons
Due to calibrated camera parameters are not absolutely accurate, the rendering results may be slightly spatially misaligned with the ground truth. In the main text, we calculate the metrics (i.e., PSNR [22], SSIM [81] and LIPIS [97]) between the aligned rendering results and the ground truth. Here, we provide the novel-view synthesis and deblurring comparisons that directly calculate metrics between the rendering results and the ground truth, as shown in Tab. A and Tab. B respectively. It shows that Deblur4DGS still outperforms state-of-the-art 4D reconstruction-based methods. More visual results for novel view synthesis in Fig. B show that our method produces more photo-realistic details. Moreover, visual results for deblurring in Fig. C show that compared with 4D reconstruction methods, Deblur4DGS produces sharper contents and fewer visual artifacts. Furthermore, the frame interpolation and video stabilization results, as well as some novel view synthesis videos are shown in the site.
D Effect of ECP and EDG
The visual results are provided in the main text. Here we provide the quantitative results in Tab. C. Without ECP and EDG, the reconstruction results are unsatisfactory due to motion blur. ECP and EDG enable processing of the camera motion blur and object motion blur respectively, thus leading to performance improvement.
E Effect of regularization terms
We validate the effect of exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$ and multi-resolution consistency regularization $\mathcal{L}_{mrc}$. In the main text, we provide the results on $288\times 512$ data. Here, we provide that on $720\times 1080$ data in Tab. D. The results are consistent for both data. $\mathcal{L}_{e}$ regularizes the object motion within the exposure time distinguished, alleviating the trivial solutions and improving performance. $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. Their combinations perform best.
F Effect of BAV canonical Gaussians
Tab. E additionally provides the quantitative results on $720\times 1080$ data, which are consistent with the ones on $288\times 512$ data. Compared with selecting a single one across the entire video (i.e., None), selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates the challenge of modeling large object motion, thus leading to performance improvement. Our method selects the canonical Gaussians corresponding to the sharper frame, which is beneficial to blur removal and performs better.
Table A: Quantitative comparisons for novel view synthesis. The metrics are directly calculated between rendering results and ground truth. The best is boldfaced, and the second is underlined.
@@ Methods
$720\times 1080$
DeformableGS [90] 16.32 / 0.644 / 0.356 16.06 / 0.711 / 0.424
4DGaussians [84] 20.58 / 0.795 / 0.186 20.31 / 0.821 / 0.258
E-D3DGS [1] 20.99 / 0.804 / 0.179 20.95 / 0.829 / 0.257
DyBluRF [72] 23.39 / 0.823 / 0.088 22.59 / 0.815 / 0.168
Shape-of-Motion [79] 26.47 / 0.908 / 0.104 25.76 / 0.896 / 0.195
Deblur4DGS (Ours) 26.69 / 0.918 / 0.052 25.93 / 0.904 / 0.102
Table B: Quantitative comparisons for deblurring. The metrics are directly calculated between rendering results and ground truth. The first category is supervised image or video deblurring methods, and the second is 4D reconstruction-based ones.
@@
$720\times 1080$
Restormer [92] 36.74 / 0.987 / 0.019 35.97 / 0.981 / 0.028
DSTNet [56] 36.49 / 0.986 / 0.015 35.24 / 0.977 / 0.027
BSSTNet [95] 36.88 / 0.988 / 0.013 35.87 / 0.983 / 0.021
DeformableGS [90] 30.97 / 0.950 / 0.096 28.95 / 0.924 / 0.189
4DGaussians [84] 31.30 / 0.951 / 0.090 29.76 / 0.929 / 0.167
E-D3DGS [1] 31.34 / 0.951 / 0.091 29.74 / 0.929 / 0.170
DyBluRF [72] 29.32 / 0.942 / 0.049 28.67 / 0.936 / 0.101
Shape-of-Motion [79] 29.51 / 0.945 / 0.095 28.35 / 0.927 / 0.187
Deblur4DGS (Ours) 31.77 / 0.972 / 0.033 30.48 / 0.956 / 0.081
Table C: Effect about the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG).
ECP EDG
@@
$720\times 1080$
$\times$ $\times$ 28.00 / 0.929 / 0.119 27.70 / 0.923 / 0.191
$\checkmark$ $\times$ 28.86 / 0.948 / 0.062 28.68 / 0.945 / 0.100
$\checkmark$ $\checkmark$ 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Table D: Ablation studies on regularization terms.
$\mathcal{L}_{e}$ $\mathcal{L}_{mfc}$ $\mathcal{L}_{mrc}$
@@
$720\times 1080$
$\times$ $\times$ $\times$ 28.69 / 0.947 / 0.063 28.70 / 0.946 / 0.099
$\times$ $\checkmark$ $\checkmark$ 28.86 / 0.949 / 0.060 28.86 / 0.947 / 0.098
$\checkmark$ $\times$ $\times$ 28.73 / 0.947 / 0.060 28.74 / 0.945 / 0.099
$\checkmark$ $\checkmark$ $\times$ 28.87 / 0.948 / 0.060 28.78 / 0.946 / 0.098
$\checkmark$ $\times$ $\checkmark$ 28.86 / 0.948 / 0.060 28.79 / 0.946 / 0.099
$\checkmark$ $\checkmark$ $\checkmark$ 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Table E: Effect of blur-aware variable (BAV) canonical Gaussians. ‘None’ denotes selecting a single one across the entire video.
@@
$720\times 1080$
None 28.67 / 0.946 / 0.065 28.66 / 0.945 / 0.104
w/o Blur-Aware 28.85 / 0.949 / 0.062 28.83 / 0.947 / 0.098
Ours 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
@@
Figure B: Visual comparisons of novel-view synthesis on the $720\times 1080$ images. Our method produces more photo-realistic details in both static and dynamic areas, as marked with yellow and red boxes respectively.

### v2

@@ Methods
None 22.13 / 0.988 / 0.017 22.30 / 0.988 / 0.016
w/o Blur-Aware 22.29 / 0.989 / 0.016 22.57 / 0.989 / 0.015
Ours 22.36 / 0.989 / 0.015 22.63 / 0.990 / 0.014
Conclusions
In this work, we propose Deblur4DGS, a 4D Gaussian Splatting framework to reconstruct a high-quality 4D model from blurry monocular video. In particular, with the explicit motion trajectory modeling, we propose to transform the challenging continuous dynamic representation estimation within an exposure time into the exposure time estimation, where a series of regularizations are suggested to tackle the under-constrained optimization. Besides, a blur-aware variable canonical Gaussians is present to represent objects with large motion better. Beyond novel-view synthesis, Deblur4DGS can improve blurry video quality from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive results show Deblur4DGS outperforms state-of-the-art 4D reconstruction methods.
@@ Conclusions
  • More comparison results in Sec.C.
  • More ablation results in Sec.D.
@@ A Structure of Camera Motion Predictor
The structure of camera motion predictor is provided in fig. 4. It first embeds the camera pose $\mathbf{P}_{t}$ to a higher dimensional space using high frequency encoder (Mildenhall et al. 2021). Then, three FC blocks are stacked, each consisting of an FC layer followed by a ReLU operation. Finally, we deploy two heads to predict the camera pose at exposure start and end (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$), respectively.
@@ B Other Regularization Terms
Following Shape-of-Motion (Wang et al. 2024), we use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, including mask regularization $\mathcal{L}_{mask}$, 2D tracks regularization $\mathcal{L}_{track}$ and distance-preserving regularization $\mathcal{L}_{rigid}$.
@@
$\hat{\mathbf{M}}_{t}^{B}(u,v)=\texttt{max}\{\hat{\mathbf{M}}_{t,1}(u,v),\hat{\mathbf{M}}_{t,2}(u,v),...,\hat{\mathbf{M}}_{t,N}(u,v)\}.$ (16)
@@
$\mathcal{L}_{mask}=\mathcal{L}_{1}(\hat{\mathbf{M}}_{t}^{B},\mathbf{M}_{t}),$ (17)
where $\mathbf{M}_{t}$ is the mask obtained by applying SAM2 (Ravi et al. 2024) to the ground truth blurry frame. Besides, we render the 2D tracks $\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}$ from a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise it by the lifted long-range 2D tracks $\mathbf{U}_{t\rightarrow t^{\prime}}$ that are extracted from TAPIR (Doersch et al. 2023), i.e.,
$\mathcal{L}_{track}=\mathcal{L}_{1}(\hat{\mathbf{U}}_{t\rightarrow t^{\prime}},\mathbf{U}_{t\rightarrow t^{\prime}}).$ (18)
@@
[…]
@@
Figure 4: Structure of camera motion predictor.
@@
$\mathcal{L}_{oth}=\lambda_{mask}\mathcal{L}_{mask}+\lambda_{track}\mathcal{L}_{track}+\lambda_{rigid}\mathcal{L}_{rigid}.$ (20)
@@
Figure 5: Effect of regularization terms $\mathcal{L}_{reg}$. $\mathcal{L}_{reg}$ includes exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$, and multi-resolution consistency regularization $\mathcal{L}_{mrc}$.
Figure 6: Some examples of our captured real-world blurry videos by the Redmi K50 Ultra smartphone.
[…]
Figure 7: Effect of BAV canonical Gaussians.
C More Comparison Results
fig. 6 shows some examples of real-world videos captured by the Redmi K50 Ultra smartphone. We evaluate methods on four tasks, i.e., novel-view synthesis, deblurring, frame interpolation, and video stabilization. Per-scene results for novel-view synthesis are summarized in table 7, table 8, table 9 and table 10. Rendering speed comparisons are in table 11. More visual comparisons for novel-view synthesis are in fig. 8 and fig. 9. They show that our Deblur4DGS produces more photo-realistic details and fewer visual artifacts. Besides, table 12, table 14 and table 16 summarize the results for deblurring, frame interpolation, and video stabilization on synthetic videos. table 13, table 15, and table 17 summarize those on real-world videos. Deblur4DGS outperforms state-of-the-art 4D reconstruction methods and has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner. Furthermore, visual results for deblurring in fig. 10 show that compared with 4D reconstruction methods, Deblur4DGS produces sharper contents. We provide some videos in the local HTML file.
D More Ablation Results
Visual results in fig. 5 validate the effectiveness of the proposed regularization terms. Without these regularization terms, notable artifacts appear in dynamic areas. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. In addition, visual results in fig. 7 validate the effectiveness of blur-aware variable (BAV) canonical Gaussians.
Table 7: Per-Scene results for novel view synthesis on $288\times 512$ synthetic data.
@@ Methods
Seesaw
[…]
Street
[…]
Basketball
[…]
Children
[…]
Sailor
[…]
Third
[…]
Man
[…]
Women
DeformableGS 16.74/0.557/0.482 18.27/0.695/0.237 12.13/0.529/0.587 15.34/0.645/0.306 15.04/0.598/0.472 12.89/0.491/0.539 16.91/0.725/0.284 19.48/0.747/0.273 14.66/0.618/0.257
4DGaussians 24.31/0.870/0.130 24.42/0.887/0.104 17.58/0.605/0.326 25.32/0.897/0.117 18.96/0.691/0.344 18.98/0.722/0.323 27.47/0.937/0.089 24.49/0.872/0.136 16.32/0.735/0.209
E-D3DGS 25.59/0.893/0.111 25.16/0.893/0.111 18.37/0.677/0.275 24.71/0.882/0.129 22.34/0.809/0.255 19.91/0.747/0.300 28.56/0.939/0.087 24.52/0.873/0.137 18.67/0.755/0.185
Shape-of-Motion 30.95/0.959/0.088 26.91/0.929/0.086 28.56/0.947/0.073 23.31/0.865/0.180 23.31/0.841/0.306 22.07/0.894/0.190 29.89/0.960/0.074 25.82/0.947/0.147 23.72/0.851/0.151
SplineGS 28.55/0.948/0.095 28.48/0.953/0.075 24.55/0.924/0.082 26.02/0.915/0.129 22.21/0.777/0.438 25.31/0.876/0.232 29.65/0.960/0.076 25.89/0.908/0.149 23.83/0.848/0.129
DyBluRF 29.54/0.935/0.076 26.46/0.936/0.080 26.40/0.938/0.085 26.24/0.907/0.059 25.32/0.903/0.111 25.27/0.905/0.123 28.57/0.937/0.070 24.40/0.926/0.103 22.23/0.860/0.105
BARD-GS 29.78/0.940/0.060 27.12/0.944/0.057 28.25/0.939/0.067 26.25/0.923/0.083 25.67/0.910/0.109 25.21/0.901/0.111 29.56/0.961/0.051 26.87/0.934/0.067 23.56/0.858/0.087
Deblur4DGS (Ours) 30.84/0.965/0.047 28.45/0.954/0.049 29.01/0.953/0.048 27.00/0.932/0.065 26.05/0.916/0.087 25.72/0.910/0.097 29.82/0.964/0.047 28.53/0.957/0.032 23.55/0.865/0.066
Table 8: Per-Scene results for novel view synthesis on $720\times 1280$ synthetic data.
@@
Seesaw
@@
Basketball
@@
Sailor
[…]
Third
[…]
Man
[…]
Women
DeformableGS 16.58/0.708/0.482 17.36/0.739/0.332 11.53/0.605/0.673 15.34/0.730/0.324 16.15/0.744/0.587 12.36/0.612/0.588 17.46/0.786/0.345 19.50/0.782/0.323 14.22/0.646/0.369
4DGaussians 23.75/0.880/0.203 23.84/0.886/0.190 17.99/0.740/0.344 20.66/0.827/0.193 22.32/0.830/0.398 18.89/0.784/0.382 26.99/0.930/0.168 23.62/0.852/0.210 17.16/0.752/0.286
E-D3DGS 25.09/0.897/0.186 24.32/0.891/0.183 18.04/0.764/0.353 23.87/0.873/0.172 22.27/0.833/0.394 19.05/0.791/0.376 27.37/0.927/0.181 23.71/0.853/0.208 18.43/0.771/0.276
Shape-of-Motion 30.11/0.946/0.169 25.57/0.913/0.200 27.87/0.939/0.139 23.21/0.876/0.224 24.12/0.860/0.488 23.76/0.864/0.355 28.92/0.948/0.156 24.84/0.879/0.238 23.88/0.853/0.248
SplineGS 28.30/0.938/0.176 27.27/0.936/0.174 24.07/0.919/0.146 24.79/0.895/0.199 22.24/0.790/0.440 20.93/0.816/0.472 28.79/0.948/0.155 24.76/0.875/0.240 23.06/0.833/0.268
DyBluRF 29.14/0.915/0.134 26.56/0.930/0.128 26.01/0.928/0.135 25.84/0.902/0.147 25.12/0.893/0.245 24.97/0.903/0.220 27.17/0.928/0.124 24.30/0.919/0.151 22.23/0.851/0.153
BARD-GS 29.30/0.930/0.122 25.89/0.935/0.125 27.98/0.942/0.122 25.96/0.912/0.116 24.29/0.831/0.250 24.70/0.880/0.193 28.90/0.948/0.112 25.84/0.920/0.101 24.20/0.890/0.113
Deblur4DGS (Ours) 30.77/0.956/0.093 27.52/0.942/0.107 28.22/0.944/0.093 26.93/0.931/0.109 25.68/0.905/0.235 24.91/0.897/0.185 28.96/0.955/0.104 28.05/0.948/0.064 23.50/0.865/0.114
Table 9: Per-Scene results for novel view synthesis on real-world videos captured by Redmi K50 Ultra.
@@
Running
@@
Boy